# On the Classical Shadow Nonparametric Bootstrap

### v1

@@ Abstract
One can think of classical shadows as a Monte Carlo experiment involving random draws of unitary operators. Likewise, bootstrap resampling can also be viewed as a Monte Carlo experiment applied to a given data set. Therefore, bootstrapping $B$ replications of a classical shadow sample of size $N$ is leveraging that data as if we had drawn $B$ classical shadow size-$N$ samples instead of a single one. Hence, bootstrapping greatly enhances the possibilities of predicting and analyzing quantum state properties with synthetic data.
@@
Repeating these measurement $N$ times produces what is called the classical shadow of $\rho,$ namely $\{\hat{\rho}_{1},\ldots,\hat{\rho}_{N}\},$ with $\mathbb{E}_{\rho}[\hat{\rho}_{i}]$ = $\rho,$ where $\mathbb{E}_{\rho}$ is the expectation under the data generating process determined by the unknown quantum state $\rho.$ Hence, the sample average of the classical shadow is one straightforward target of interest. More generally, the targeted predictions include linear functions of the underlying density matrix $\rho,$ or even nonlinear functions of $\rho,$ such as entanglement entropy. Besides expectation values, other examples of linear functionals include the fidelity with a pure target state, the probability distribution governing the possible outcomes of a measurement, among others.
@@
Hence, the classical shadow provides an unbiased estimator of the linear functional $\mathrm{tr}(O_{i}\rho)$. To avoid outlier corruption, huang2020predicting split the classical shadow sample into $K$ equally-sized independent sub-samples of size $N$ and estimate sample means for each, yielding the following median-of-mean estimator:
@@
Figure 1: Gaussian distribution of median-of-mean estimator with $K$ = 10 for $N$ = 1000, for an observable $o_{1}$ of a 10-qubit circuit involving rotations detailed in the next section, using the estimator defined in equation (2) and its variance appearing in (4).
@@
  • For each $b$ = 1, $\ldots$ we create a sample of size $N$ (called the $b^{th}$ bootstrap sample) $\hat{\rho}_{1}^{*,b},\hat{\rho}_{2}^{*,b},\dots,\hat{\rho}_{N}^{*,b}$ which is obtained by drawing $N$ observations with replacement from the original dataset $\{\hat{\rho}_{1},\dots,\hat{\rho}_{N}\}$. Each $\hat{\rho}_{i}^{*,b}\sim\widehat{F}_{N}$ i.i.d.
@@ 10    $\{M_{N}^{*(b)}\}_{b=1}^{B}$ approximates sampling distribution of classical shadow mean estimators
We use classical shadows to estimate a set $O$ of 27 observables. $O$ contains the $XX$, YYYY, and $ZZ$ correlations for all adjacent qubits in the circuit, or put another way:
@@
We implement the classical shadow nonparametric bootstrap introduced in the previous section and display the histogram of the distribution for $B$ = 1000 bootstrap samples each involving a median-of-mean estimator with parameters $K$ = 10 and $N$ = 1000, as appearing in equation (2). [^5] All of the code used to generate the figures in this paper is publicly available at https://github.com/jackhmorgan/Classical-Shadow-Nonparametric-Bootstrap. Our implementation is heavily inspired by the pennylate classical shadows tutorial found here: https://pennylane.ai/qml/demos/tutorial_classical_shadows. For illustrative purpose, we focus again on the observable $o_{1}$ of the 10-qubit circuit (but will discuss other observables as well later). Figure 3 displays three distributions, namely: (a) the Gaussian which appeared in Figure 1 representing the asymptotic representation, (b) the histogram of the $B$ = 1000 expectation value bootstrap estimates for the median-of-means, and (c) a Kernel density estimate (KDE) smoothed version of the histogram. The bootstrap distribution, either depicted by the histogram or by KDE, is non-Gaussian. We notice (a) heavy tails, (b) some hint of asymmetry and (c) similar to the Gaussian approximation a minor small sample bias. Figure 3 tells us that aside from the modest small sample bias, present both in the Gaussian and bootstrap distributions, we have a serious problem of the approximating the tails with the Gaussian density.
To shed further light on the inadequacy of the asymptotic distribution we examine the error bounds on the estimation error suggested by huang2020predicting . They show that to achieve accuracy $\varepsilon$ and confidence $1-\delta$ for $M$ observables $\{O_{i}\}_{i=1}^{M}$, the parameters $K$ and $N$ for the median-of-mean estimator are chosen as
@@
This ensures uniform accuracy over all $M$ observables using $NK$ independent classical shadows.
Figure 4 displays the error bounds for $\varepsilon$ = $0.95$ and $0.50,$ with $N$ = 500, 1000, 1500, 2000, 2500 and 3000 and $B$ = 1000. We clearly see from the box-plotted bootstrap distributions that the (asymptotic) error bounds are not tight. Hence, we should rely on the bootstrap distibutions to assess the tail behavior. Moreover, the figure also shows that the small sample bias diminishes as we increase $N,$ as is expected.

### v2

@@ Abstract
One can think of classical shadows as a Monte Carlo experiment involving random draws of unitary operators. Likewise, bootstrap resampling can also be viewed as a Monte Carlo experiment applied to a given data set. Therefore, bootstrapping $B$ replications of a classical shadow sample of size NN is leveraging that data as if we had drawn $B$ classical shadow size-NN samples instead of a single one. Hence, bootstrapping greatly enhances the possibilities of predicting and analyzing quantum state properties with synthetic data.
@@
Repeating these measurement NN times produces what is called the classical shadow of $\rho,$ namely $\{\hat{\rho}_{1},\ldots,\hat{\rho}_{N}\},$ with $\mathbb{E}_{\rho}[\hat{\rho}_{i}]$ = $\rho,$ where $\mathbb{E}_{\rho}$ is the expectation under the data generating process determined by the unknown quantum state $\rho.$ Hence, the sample average of the classical shadow is one straightforward target of interest. More generally, the targeted predictions include linear functions of the underlying density matrix $\rho,$ or even nonlinear functions of $\rho,$ such as entanglement entropy. Besides expectation values, other examples of linear functionals include the fidelity with a pure target state, the probability distribution governing the possible outcomes of a measurement, among others.
@@
Hence, the classical shadow provides an unbiased estimator of the linear functional $\mathrm{tr}(O_{i}\rho)$. To avoid outlier corruption, huang2020predicting split the classical shadow sample into $K$ equally-sized independent sub-samples of size NN and estimate sample means for each, yielding the following median-of-mean estimator:
@@
Figure 1: Gaussian distribution of median-of-mean estimator with $K$ = 10 for NN = 1000, for an observable $o_{1}$ of a 10-qubit circuit involving rotations detailed in the next section, using the estimator defined in equation (2) and its variance appearing in (4).
@@
  • For each $b$ = 1, $\ldots$ we create a sample of size NN (called the $b^{th}$ bootstrap sample) $\hat{\rho}_{1}^{*,b},\hat{\rho}_{2}^{*,b},\dots,\hat{\rho}_{N}^{*,b}$ which is obtained by drawing NN observations with replacement from the original dataset $\{\hat{\rho}_{1},\dots,\hat{\rho}_{N}\}$. Each $\hat{\rho}_{i}^{*,b}\sim\widehat{F}_{N}$ i.i.d.
@@ 10    $\{M_{N}^{*(b)}\}_{b=1}^{B}$ approximates sampling distribution of classical shadow mean estimators
We use classical shadows to estimate a set $O$ of 27 observables. $O$ contains the XXXX, YYYY, and ZZZZ correlations for all adjacent qubits in the circuit, or put another way:
@@
We implement the classical shadow nonparametric bootstrap introduced in the previous section and display the histogram of the distribution for $B$ = 1000 bootstrap samples each involving a median-of-mean estimator with parameters $K$ = 10 and NN = 1000, as appearing in equation (2). [^5] All of the code used to generate the figures in this paper is publicly available at https://github.com/jackhmorgan/Classical-Shadow-Nonparametric-Bootstrap. Our implementation is heavily inspired by the pennylate classical shadows tutorial found here: https://pennylane.ai/qml/demos/tutorial_classical_shadows. For illustrative purpose, we focus again on the observable $o_{1}$ of the 10-qubit circuit (but will discuss other observables as well later). Figure 3 displays three distributions, namely: (a) the Gaussian which appeared in Figure 1 representing the asymptotic representation, (b) the histogram of the $B$ = 1000 expectation value bootstrap estimates for the median-of-means, and (c) a Kernel density estimate (KDE) smoothed version of the histogram. The bootstrap distribution, either depicted by the histogram or by KDE, is non-Gaussian. We notice (a) heavy tails, (b) some hint of asymmetry and (c) similar to the Gaussian approximation a minor small sample bias. Figure 3 tells us that aside from the modest small sample bias, present both in the Gaussian and bootstrap distributions, we have a serious problem of the approximating the tails with the Gaussian density.
To shed further light on the inadequacy of the asymptotic distribution we examine the error bounds on the estimation error suggested by huang2020predicting . They show that to achieve accuracy $\varepsilon$ and confidence $1-\delta$ for $M$ observables $\{O_{i}\}_{i=1}^{M}$, the parameters $K$ and NN for the median-of-mean estimator are chosen as
@@
This ensures uniform accuracy over all $M$ observables using NKNK independent classical shadows.
Figure 4 displays the error bounds for $\varepsilon$ = $0.95$ and $0.50,$ with NN = 500, 1000, 1500, 2000, 2500 and 3000 and $B$ = 1000. We clearly see from the box-plotted bootstrap distributions that the (asymptotic) error bounds are not tight. Hence, we should rely on the bootstrap distibutions to assess the tail behavior. Moreover, the figure also shows that the small sample bias diminishes as we increase $N,$ as is expected.